At a certain baseball park the reserved seats cost $1 and the other seats were 50 cents each. If 3600 tickets were sold in all and the total amount received was $2225, how many tickets of each kind were sold?
Answer by mananth(16946)

x= number of $1 tickets
y= number of $0.5 tickets

1.00 x + 1.00 y = 3600.00 .............1
Total value
1.00 x 0.50 y = 2225.00 .............2
Eliminate y
multiply (1)by -0.50
Multiply (2) by 1.00
-0.50 x -0.50 y = -1800.00
1.00 x + 0.50 y = 2225.00
Add the two equations
0.50 x = 425.00
/ 0.50
x = 850.00
plug value of x in (1)
1.00 x + 1.00 y = 3600.00
850.00 + y = 3600.00
y = 3600.00 -850.00
y = 2750.00
y = 2750.00
x= 850 number of $1 tickets
y= 2750 number of $0.5 tickets
m.ananth@hotmail.ca